Solve 3(x-5)+3x=4x+5
Answer by tutorcecilia(2144)

3(x-5)+3x=4x+5
3x-15+3x=4x+5
6x-15=4x+5
6x-4x-15=4x-4x+5
2x-15=5
2x-15+15=5+15
2x=20
2x/2=20/2
x=10
.
check by plugging (x=10) back into the original equation and solve:
3(10-5)+3x=4x+5
3(5)+3(10)=4(10)+5
15+30=40+5
45=45 [checks out]